# Universality of Layer-Level Entropy-Weighted Quantization Beyond Model Architecture and Size

### v1

@@ 1 Introduction
The widespread adoption of LLMs has been constrained by their substantial computational and memory requirements, particularly as model sizes continue to grow exponentially (Brown et al., 2020; Chowdhery et al., 2022; Hoffmann et al., 2022). As LLMs become increasingly integral to various applications, from natural language processing to automated reasoning, the need for efficient deployment solutions becomes more critical. A typical LLM with 7-70B parameters requires 14-140GB of memory in full precision, making deployment challenging even in well-resourced data centers and practically impossible on edge devices or consumer hardware.
@@
Neural network quantization has emerged as a promising approach to address these deployment challenges by reducing resource usage and accelerating inference through lowering the bit precision of model parameters (Choi et al., 2018; Hubara et al., 2021; Yao et al., 2022; Park et al., 2022; Gholami et al., 2022). We can define quantization as a way of reducing data precision from typical 32-bit or bfloat16 to smaller bits (<=<=< = 8-bit), which in turn lowers the size of the model and speeds up matrix multiplications involved in the attention mechanism. While quantization offers significant memory savings—potentially reducing model size by 75% or more—maintaining model performance under reduced precision remains a fundamental challenge.
@@
Quantization techniques can be broadly categorized into uniform precision quantization and mixed precision quantization. While uniform precision quantization is widely applied to reduce the size of transformer layers in LLMs, its indiscriminate application often leads to significant performance degradation. This degradation occurs because different layers in transformer-based models exhibit varying sensitivity to quantization, necessitating more nuanced approaches (Wang et al., 2018; Cai and Vasconcelos, 2020; Zadeh and Moshovos, 2020; Ganesh et al., 2021). For example, early layers processing raw input tokens and final layers producing output logits typically require higher precision than intermediate layers.
@@
Recent research has focused on addressing the challenges posed by outlier activations, which represent a key impediment to effective uniform low-precision quantization. Mixed-precision quantization has shown promise in mitigating this issue by maintaining outlier channels at higher precision (Dettmers et al., 2022a; Zhao et al., 2024; Ashkboos et al., 2023; Zhao et al., 2024). Another emerging approach is invariant random rotation, which aims to suppress outliers and enable more effective uniform low-precision quantization (Ashkboos et al., 2024; Liu et al., 2024; Wei et al., 2023). While both methods improve the signal-to-quantization noise ratio and reduce quantization errors locally, they have yet to demonstrate substantial performance advantages over 16-bit precision models. For instance, SpinQuant (Liu et al., 2024) applied at 4-bit precision to Llama-3-8b (Meta, 2024) shows approximately 20% higher perplexity compared to the 16-bit baseline, despite significant optimization efforts.
@@
Alternative approaches such as SmoothQuant (Xiao et al., 2024) and Activation-Aware Quantization (AWQ) (Yuan et al., 2023; Lin et al., 2024) have been developed to enable effective 8-bit quantization of both weights and activations. These methods employ sophisticated techniques, including offline migration of quantization difficulty from activations to weights and channel equalization. However, they typically require access to the entire model for activation distribution analysis, making them impractical in resource-constrained environments (Kim et al., 2023a). This is why weight-only quantization represents a more suitable use case for model compression, exemplified by methods like GPTQ (Frantar et al., 2022) or FineQuant (Kim et al., 2023c). GPTQ converts quantized weights to float16 during inference for matrix multiplication operations and FineQuant uses a fine-grained quantization algorithm that incorporates group-wise quantization and adaptive selection of granularity. While these approaches can achieve performance gains through reduced data loading with minimal accuracy loss, they present practical limitations in production environments specially in distributed settings. For example, handling long context lengths and batch processing increases the memory footprint of key-value (KV) cache substantially. Or uniform quantization of all transformer blocks happen to increase perplexity significantly. Still, even with these challenges the weight-only quantization seems to be a promising approach for the purposes of this paper. So we would like to pose the following important question.
@@
To provide a production-grade answer, we introduce Entropy-Weighted Quantization (EWQ), a systematic framework for selective model compression that preserves performance while substantially reducing memory requirements. Our approach extends the theoretical foundations of information-theoretic neural network compression (Park et al., 2017; Xu et al., 2018; Dong et al., 2019c) to address the unique challenges of large language models. While entropy-based quantization has proven effective for traditional machine learning architectures (Park et al., 2017) and has recently shown promise in quantization-aware training of LLMs (Shen et al., 2024), its application to runtime weight-only quantization remains unexplored. This gap is particularly significant given that LLMs exhibit markedly heterogeneous entropy distributions across their transformer layers—a characteristic that distinguishes them from smaller neural architectures and necessitates more sophisticated quantization strategies.
Our approach’s distinguishing feature is its ability to facilitate efficient on-the-fly quantization, making it particularly well-suited for deployment on consumer-grade hardware with limited resources. Unlike traditional mixed-precision methods that demand significant computational overhead for activation analysis (Dettmers et al., 2022a; Kim et al., 2023b), EWQ’s weight-centric approach allows for rapid deployment while maintaining adaptability to different hardware constraints. By intelligently mapping the entropy characteristics of LLM weights to appropriate precision levels, EWQ achieves an optimal balance between model performance and resource efficiency. This ensures that quantized models maintain competitive performance with their full-precision counterparts while significantly reducing memory footprint and accelerating inference speed. The architecture-agnostic nature of EWQ is further demonstrated by the emergence of a universal approximator for the quantization of transformer layers, enabling rapid quantization—termed FastEWQ—without the need to load weights.
@@ 2.1 Model Quantization
In more technical terms, model (neural network) quantization refers to the process of reducing the numerical precision of weights and activations from 32-bit floating point to lower bit-width representations (typically 8-bit or 4-bit), quantization achieves significant reductions in memory footprint and computational requirements while preserving model functionality (Choi et al., 2018; Hubara et al., 2021). Modern quantization approaches can be broadly categorized into two paradigms:
  • Uniform quantization: Applies identical precision reduction across all model components, enabling straightforward implementation but often resulting in significant accuracy degradation for sensitive layers (Yao et al., 2022).
  • Mixed-precision quantization: Allocates higher precision to critical layers identified through sensitivity analysis, achieving better accuracy preservation at the cost of increased implementation complexity (Dettmers et al., 2022a; Zhao et al., 2024).
Recent advances in post-training quantization (PTQ) have demonstrated particular promise for LLM deployment. Dettmers et al. (2022a) introduced layer-wise adaptive mixed precision for GPT-3 models, maintaining 16-bit precision only for outlier-dominated attention heads. Frantar et al. (2022) developed a second-order quantization approach that minimizes layer-wise reconstruction errors, enabling 4-bit quantization of LLaMA models with minimal accuracy loss. The BitLinear layer proposed by Ashkboos et al. (2023) achieves extreme 1.58-bit quantization through entropy-driven logarithmic representations, though with increased computational overhead.
@@ 2.2 Block Sensitivity Analysis
The transformer architecture’s layered structure exhibits significant heterogeneity in quantization sensitivity across blocks. Early work by Devlin et al. (2019) demonstrated that initial encoder layers in BERT models capture fundamental syntactic features highly sensitive to precision reduction, while deeper layers encode semantic relationships more tolerant of quantization. This phenomenon was formalized by Dong et al. (2019a, b); Shen et al. (2019) through Hessian-based sensitivity analysis, establishing that attention blocks typically require 2-4× higher precision than feedforward layers.
@@
  • Progressive quantization: Gradually increases quantization intensity from output to input layers, preserving early layer precision (Zadeh and Moshovos, 2020).
  • Attention-aware allocation: Assigns higher precision to query/key matrices than value projections to maintain attention fidelity (Passban et al., 2021).
  • Task-adaptive thresholds: Dynamically adjusts layer precision based on downstream task gradients (Kim et al., 2023a).
In particular, Yuan et al. (2023) introduced activation-aware singular value decomposition (ASVD), a post-training compression that addresses challenges in low-rank factorization by managing activation outliers through the scaling of the weight matrix based on the activation distributions, enhancing the accuracy of decomposition. Additionally, it employs an iterative calibration process to optimize layer-specific decomposition, considering the varying sensitivity of different LLM layers. Experiments demonstrate that ASVD can compress networks by 10-20% without compromising performance. Furthermore, by applying low-rank decomposition to projection matrices in the self-attention module, ASVD achieves up to 50% reduction in KV cache memory requirements without performance degradation.
@@ 2.3 Information-Theoretic Approaches
The relationship between parameter entropy and quantization robustness originates from fundamental rate-distortion theory (Cover and Thomas, 2006), where entropy establishes theoretical bounds on lossy compression. In deep learning, Park et al. (2017) first operationalized this connection by demonstrating that weight matrices with Shannon entropy $H(W)\leq 4$ bits/parameter could withstand 4-bit quantization with less than 1% accuracy drop. Their analysis revealed that entropy correlates with parameter redundancy—layers learning simpler patterns (e.g., smooth feature detectors) naturally exhibit lower entropy and greater quantization tolerance (Jin et al., 2021).
@@
Calculating the full entropy for all $n$ transformer layers leads to a complexity of $O(n)$, which becomes prohibitive in billion-parameter models. We mitigate this by focusing solely on weight-only quantization, maintaining $O(n)$ time and space complexity. Another challenge is architecture-specific sensitivity. Existing thresholds, such as Park’s 4-bit boundary (Park et al., 2017), fail for heterogeneous architectures like Mixture of Experts (MoE) models. Instead, we derive architecture-agnostic criteria using FastEWQ to generalize across multiple model families.
A crucial step for any post-training quantization is downloading the model weights, which significantly limits access to hard disk resources. Static quantization policies (Ashkboos et al., 2023) cannot adapt to varying cluster resources. Our on-the-fly optimization framework operates in $O(n)$ time per resource update for EWQ and $O(1)$ time with FastEWQ, maintaining Pareto-optimal accuracy-efficiency trade-offs (Abdolrashidi et al., 2021). This synthesis enables the first information-theoretic quantization system that simultaneously achieves sublinear time entropy estimation, cross-architecture validity, and real-time adaptation to deployment constraints.
@@ 2.5 The Case for Mixed Quantization
Fully 4-bit Quantization <10% <10% Poor performance on both metrics; unsuitable for high-precision tasks.
@@
In the context of QA tasks, the ResQ method applies mixed-precision quantization to LLMs, demonstrating that such approaches can maintain model performance while reducing computational costs (Saxena et al., 2024). Similarly, the SliM-LLM framework employs a salience-driven mixed-precision quantization scheme, achieving efficient LLMs with high accuracy (Huang et al., 2025). These findings align with our later observations that mixed-precision quantization strategies can effectively balance model efficiency and performance in QA tasks.
@@ 3.4 Optimized Distribution of LLM transformer Blocks in Deployment Clusters
For deployment scenarios with severe resource constraints, such as edge devices or mobile phones, we can adapt this methodology to employ a 4-3bit (or even 2-bit (Chee et al., 2024)). combination instead of the standard 8-4bit approach. In this configuration, high-entropy blocks are preserved at 4-bit precision while lower-entropy blocks are further compressed to 3-bit precision. Our experiments with this configuration demonstrate that for models deployed on devices with less than 2GB of available RAM, the 4-3bit combination can reduce the model footprint by an additional 18-25% compared to uniform 4-bit quantization while maintaining acceptable accuracy degradation of less than 5% on standard benchmarks. Furthermore, the block distribution algorithm dynamically adjusts to network topology, prioritizing block placement that minimizes cross-machine communication during inference. This is particularly important for deployment clusters with heterogeneous hardware, as the algorithm can assign computation-intensive blocks to machines with better processing capabilities while memory-intensive operations can be directed to machines with larger available memory. The result is a holistic optimization that considers not just quantization but also the operational characteristics of the deployment environment, leading to more efficient resource utilization and reduced inference latency.
@@ 4.3 Feature Importance Analysis
The prominence of exec_index can be attributed to the inherent processing structure of transformer models. Early layers primarily capture local syntactic features, while deeper layers encode more abstract semantic representations. Quantizing blocks inappropriately across this hierarchy can lead to a degradation in model performance, as different layers contribute variably to the model’s overall function. This aligns with analyses that highlight the distinct roles of transformer layers in processing information (Kobayashi et al., 2024).
@@
  • Redundancy Scaling: Larger blocks, such as feed-forward networks (FFNs), often exhibit higher parameter redundancy, making them more suitable candidates for quantization. Research indicates that despite their substantial parameter count, FFNs can be compressed with minimal impact on performance (Pires et al., 2023).
@@
A possible reason for the predictive power of the execution index might stem from information bottleneck principles (Tishby and Zaslavsky, 2015). As information propagates through the transformer, we have
$I(Y;T_{i})=I(X;T_{i})-I(X;T_{i}|Y)$ (1)
where $T_{i}$ represents the $i^{th}$ block, $X$ the input, and $Y$ the target. Middle blocks optimize the trade-off between input compression $I(X;T_{i})$ and predictive relevance $I(Y;T_{i})$, leading to maximally compressed (low entropy) representations (Shwartz-Ziv and Tishby, 2017). This compression manifests as parameter redundancy, creating natural quantization opportunities.

### v2

@@ 1 Introduction
The widespread adoption of LLMs has been constrained by their substantial computational and memory requirements, particularly as model sizes continue to grow exponentially [Brown et al., 2020, Chowdhery et al., 2022, Hoffmann et al., 2022]. As LLMs become increasingly integral to various applications, from natural language processing to automated reasoning, the need for efficient deployment solutions becomes more critical. A typical LLM with 7-70B parameters requires 14-140GB of memory in full precision, making deployment challenging even in well-resourced data centers and practically impossible on edge devices or consumer hardware.
@@
Neural network quantization has emerged as a promising approach to address these deployment challenges by reducing resource usage and accelerating inference through lowering the bit precision of model parameters [Choi et al., 2018, Hubara et al., 2021, Yao et al., 2022, Park et al., 2022, Gholami et al., 2022]. We can define quantization as a way of reducing data precision from typical 32-bit or bfloat16 to smaller bits (<=<=< = 8-bit), which in turn lowers the size of the model and speeds up matrix multiplications involved in the attention mechanism. While quantization offers significant memory savings—potentially reducing model size by 75% or more—maintaining model performance under reduced precision remains a fundamental challenge.
@@
Quantization techniques can be broadly categorized into uniform precision quantization and mixed precision quantization. While uniform precision quantization is widely applied to reduce the size of transformer layers in LLMs, its indiscriminate application often leads to significant performance degradation. This degradation occurs because different layers in transformer-based models exhibit varying sensitivity to quantization, necessitating more nuanced approaches [Wang et al., 2018, Cai and Vasconcelos, 2020, Zadeh and Moshovos, 2020, Ganesh et al., 2021]. For example, early layers processing raw input tokens and final layers producing output logits typically require higher precision than intermediate layers.
@@
Recent research has focused on addressing the challenges posed by outlier activations, which represent a key impediment to effective uniform low-precision quantization. Mixed-precision quantization has shown promise in mitigating this issue by maintaining outlier channels at higher precision [Dettmers et al., 2022a, Zhao et al., 2024, Ashkboos et al., 2023, Zhao et al., 2024]. Another emerging approach is invariant random rotation, which aims to suppress outliers and enable more effective uniform low-precision quantization [Ashkboos et al., 2024, Liu et al., 2024, Wei et al., 2023]. While both methods improve the signal-to-quantization noise ratio and reduce quantization errors locally, they have yet to demonstrate substantial performance advantages over 16-bit precision models. For instance, SpinQuant [Liu et al., 2024] applied at 4-bit precision to Llama-3-8b [Meta, 2024] shows approximately 20% higher perplexity compared to the 16-bit baseline, despite significant optimization efforts.
@@
Alternative approaches such as SmoothQuant [Xiao et al., 2024] and Activation-Aware Quantization (AWQ) [Yuan et al., 2023, Lin et al., 2024] have been developed to enable effective 8-bit quantization of both weights and activations. These methods employ sophisticated techniques, including offline migration of quantization difficulty from activations to weights and channel equalization. However, they typically require access to the entire model for activation distribution analysis, making them impractical in resource-constrained environments [Kim et al., 2023a]. This is why weight-only quantization represents a more suitable use case for model compression, exemplified by methods like GPTQ [Frantar et al., 2022] or FineQuant [Kim et al., 2023c]. GPTQ converts quantized weights to float16 during inference for matrix multiplication operations and FineQuant uses a fine-grained quantization algorithm that incorporates group-wise quantization and adaptive selection of granularity. While these approaches can achieve performance gains through reduced data loading with minimal accuracy loss, they present practical limitations in production environments specially in distributed settings. For example, handling long context lengths and batch processing increases the memory footprint of key-value (KV) cache substantially. Or uniform quantization of all transformer blocks happen to increase perplexity significantly. Still, even with these challenges the weight-only quantization seems to be a promising approach for the purposes of this paper. So we would like to pose the following important question.
@@
To provide a production-grade answer, we introduce Entropy-Weighted Quantization (EWQ), a systematic framework for selective model compression that preserves performance while substantially reducing memory requirements. Our approach extends the theoretical foundations of information-theoretic neural network compression [Park et al., 2017, Xu et al., 2018, Dong et al., 2019c] to address the unique challenges of large language models. While entropy-based quantization has proven effective for traditional machine learning architectures [Park et al., 2017] and has recently shown promise in quantization-aware training of LLMs [Shen et al., 2024], its application to runtime weight-only quantization remains unexplored. This gap is particularly significant given that LLMs exhibit markedly heterogeneous entropy distributions across their transformer layers—a characteristic that distinguishes them from smaller neural architectures and necessitates more sophisticated quantization strategies.
Our approach’s distinguishing feature is its ability to facilitate efficient on-the-fly quantization, making it particularly well-suited for deployment on consumer-grade hardware with limited resources. Unlike traditional mixed-precision methods that demand significant computational overhead for activation analysis [Dettmers et al., 2022a, Kim et al., 2023b], EWQ’s weight-centric approach allows for rapid deployment while maintaining adaptability to different hardware constraints. By intelligently mapping the entropy characteristics of LLM weights to appropriate precision levels, EWQ achieves an optimal balance between model performance and resource efficiency. This ensures that quantized models maintain competitive performance with their full-precision counterparts while significantly reducing memory footprint and accelerating inference speed. The architecture-agnostic nature of EWQ is further demonstrated by the emergence of a universal approximator for the quantization of transformer layers, enabling rapid quantization—termed FastEWQ—without the need to load weights.
@@ 2.1 Model Quantization
In more technical terms, model (neural network) quantization refers to the process of reducing the numerical precision of weights and activations from 32-bit floating point to lower bit-width representations (typically 8-bit or 4-bit), quantization achieves significant reductions in memory footprint and computational requirements while preserving model functionality [Choi et al., 2018, Hubara et al., 2021]. Modern quantization approaches can be broadly categorized into two paradigms:
  • Uniform quantization: Applies identical precision reduction across all model components, enabling straightforward implementation but often resulting in significant accuracy degradation for sensitive layers [Yao et al., 2022].
  • Mixed-precision quantization: Allocates higher precision to critical layers identified through sensitivity analysis, achieving better accuracy preservation at the cost of increased implementation complexity [Dettmers et al., 2022a, Zhao et al., 2024].
Recent advances in post-training quantization (PTQ) have demonstrated particular promise for LLM deployment. Dettmers et al. [2022a] introduced layer-wise adaptive mixed precision for GPT-3 models, maintaining 16-bit precision only for outlier-dominated attention heads. Frantar et al. [2022] developed a second-order quantization approach that minimizes layer-wise reconstruction errors, enabling 4-bit quantization of LLaMA models with minimal accuracy loss. The BitLinear layer proposed by Ashkboos et al. [2023] achieves extreme 1.58-bit quantization through entropy-driven logarithmic representations, though with increased computational overhead.
@@ 2.2 Block Sensitivity Analysis
The transformer architecture’s layered structure exhibits significant heterogeneity in quantization sensitivity across blocks. Early work by Devlin et al. [2019] demonstrated that initial encoder layers in BERT models capture fundamental syntactic features highly sensitive to precision reduction, while deeper layers encode semantic relationships more tolerant of quantization. This phenomenon was formalized by Dong et al. [2019a, b], Shen et al. [2019] through Hessian-based sensitivity analysis, establishing that attention blocks typically require 2-4× higher precision than feedforward layers.
@@
  • Progressive quantization: Gradually increases quantization intensity from output to input layers, preserving early layer precision [Zadeh and Moshovos, 2020].
  • Attention-aware allocation: Assigns higher precision to query/key matrices than value projections to maintain attention fidelity [Passban et al., 2021].
  • Task-adaptive thresholds: Dynamically adjusts layer precision based on downstream task gradients [Kim et al., 2023a].
In particular, Yuan et al. [2023] introduced activation-aware singular value decomposition (ASVD), a post-training compression that addresses challenges in low-rank factorization by managing activation outliers through the scaling of the weight matrix based on the activation distributions, enhancing the accuracy of decomposition. Additionally, it employs an iterative calibration process to optimize layer-specific decomposition, considering the varying sensitivity of different LLM layers. Experiments demonstrate that ASVD can compress networks by 10-20% without compromising performance. Furthermore, by applying low-rank decomposition to projection matrices in the self-attention module, ASVD achieves up to 50% reduction in KV cache memory requirements without performance degradation.
@@ 2.3 Information-Theoretic Approaches
The relationship between parameter entropy and quantization robustness originates from fundamental rate-distortion theory [Cover and Thomas, 2006], where entropy establishes theoretical bounds on lossy compression. In deep learning, Park et al. [2017] first operationalized this connection by demonstrating that weight matrices with Shannon entropy $H(W)\leq 4$ bits/parameter could withstand 4-bit quantization with less than 1% accuracy drop. Their analysis revealed that entropy correlates with parameter redundancy—layers learning simpler patterns (e.g., smooth feature detectors) naturally exhibit lower entropy and greater quantization tolerance [Jin et al., 2021].
@@
Calculating the complete entropy data for all $n$ transformer layers leads to a complexity of $O(n)$ including the entropy distribution for activations [Shen et al., 2024], which becomes prohibitive in billion-parameter models. We mitigate this by focusing solely on weight-only quantization, maintaining $O(n)$ time and space complexity. Another challenge is architecture-specific sensitivity. Existing thresholds, such as Park’s 4-bit boundary [Park et al., 2017], fail for heterogeneous architectures like Mixture of Experts (MoE) models. Instead, we derive architecture-agnostic criteria using FastEWQ to generalize across multiple model families.
A crucial step for any post-training quantization is downloading the model weights, which significantly limits access to hard disk resources. Static quantization policies [Ashkboos et al., 2023] cannot adapt to varying cluster resources. Our on-the-fly optimization framework operates in $O(n)$ time per resource update for EWQ and $O(1)$ time with FastEWQ, maintaining Pareto-optimal accuracy-efficiency trade-offs [Abdolrashidi et al., 2021]. This synthesis enables the first information-theoretic quantization system that simultaneously achieves sublinear time entropy estimation, cross-architecture validity, and real-time adaptation to deployment constraints.
@@ 2.5 The Case for Mixed Quantization
Fully 4-bit Quantization <35% <12% Poor performance on both metrics; unsuitable for high-precision tasks.
@@
In the context of QA tasks, the ResQ method applies mixed-precision quantization to LLMs, demonstrating that such approaches can maintain model performance while reducing computational costs [Saxena et al., 2024]. Similarly, the SliM-LLM framework employs a salience-driven mixed-precision quantization scheme, achieving efficient LLMs with high accuracy [Huang et al., 2025]. These findings align with our later observations that mixed-precision quantization strategies can effectively balance model efficiency and performance in QA tasks.
@@ 3.4 Optimized Distribution of LLM transformer Blocks in Deployment Clusters
For deployment scenarios with severe resource constraints, such as edge devices or mobile phones, we can adapt this methodology to employ a 4-3bit (or even 2-bit [Chee et al., 2024]). combination instead of the standard 8-4bit approach. In this configuration, high-entropy blocks are preserved at 4-bit precision while lower-entropy blocks are further compressed to 3-bit precision. Our experiments with this configuration demonstrate that for models deployed on devices with less than 2GB of available RAM, the 4-3bit combination can reduce the model footprint by an additional 18-25% compared to uniform 4-bit quantization while maintaining acceptable accuracy degradation of less than 5% on standard benchmarks. Furthermore, the block distribution algorithm dynamically adjusts to network topology, prioritizing block placement that minimizes cross-machine communication during inference. This is particularly important for deployment clusters with heterogeneous hardware, as the algorithm can assign computation-intensive blocks to machines with better processing capabilities while memory-intensive operations can be directed to machines with larger available memory. The result is a holistic optimization that considers not just quantization but also the operational characteristics of the deployment environment, leading to more efficient resource utilization and reduced inference latency.
@@ 4.3 Feature Importance Analysis
The prominence of exec_index can be attributed to the inherent processing structure of transformer models. Early layers primarily capture local syntactic features, while deeper layers encode more abstract semantic representations. Quantizing blocks inappropriately across this hierarchy can lead to a degradation in model performance, as different layers contribute variably to the model’s overall function. This aligns with analyses that highlight the distinct roles of transformer layers in processing information [Kobayashi et al., 2024].
@@
  • Redundancy Scaling: Larger blocks, such as feed-forward networks (FFNs), often exhibit higher parameter redundancy, making them more suitable candidates for quantization. Research indicates that despite their substantial parameter count, FFNs can be compressed with minimal impact on performance [Pires et al., 2023].
@@
A possible explanation for the predictive power of the execution index may lie in the principles of the information bottleneck [Tishby and Zaslavsky, 2015]. Assuming that the transformer blocks $T_{i}$ form a Markov chain [Makkuva et al., 2024], the propagation of information through the transformer can be described by the mutual information
$I(T_{i};T_{i+1})=H(T_{i})-H(T_{i}|T_{i+1})$
where $H(T_{i})$ denotes the entropy of the $i^{th}$ block, and $H(T_{i}|T_{i+1})$ represents the conditional entropy of $T_{i}$ given $T_{i+1}$. Mutual information quantifies the average number of relevant bits that $T_{i}$ retains about $T_{i+1}$ [Shwartz-Ziv and Tishby, 2017]. Since mutual information is bounded below by zero, the inequality $H(T_{i})\geq H(T_{i}|T_{i+1})$ holds, implying that lower entropy in block $T_{i}$ leads to a tighter bound on the conditional entropy. Consequently, this constrains the mutual information to smaller values. Thus, it follows naturally that the execution index plays a crucial role in optimal quantization through the lens of mutual information.